OpenStudy (anonymous):

1. Solve the equation. Round the answer to the nearest tenth. 3.2x = 4.6 (Points : 1)

OpenStudy (pdoctor):

Divide both sides by 3.2 \[(3.2x)/3.2=4.6/3.2\] Therefore, x= 4.6/3.2 x = 1.4375 x = 1.4

OpenStudy (anonymous):